Prove that the sum of AP is sn equals(2a+(n-1)d)
Answer by stanbon(75887)

S(n) = ....... a + (a+d) + (a+2d)+......+(a+(n-1)d)
-----
S(n) = (a+(n-1)d) + (a+(n-2)d) + (a+(n-3)d)+....+ a
----
Both lines are the sum of n terms of the same AP.
----
Add the 2 lines to get::
2*S(n) = (2a+(n-1)d) + (2a+(n-1)d)+....+(2a+(n-1)d)
-----
2*S(n) = n(2a+(n-1)d)
-----
Then S(n) = (n/2)(2a+(n-1)d) = (n/2)(a + [a+(n-1)d])
=========================================================
Cheers,
Stan H.
=====================